A rocket beigns from rest and accelerates upward at +0.5m/s^2.
a) At what time will the rocket reach a height of 1000 m?
b) What will be the rocket's velocity at that point?
**BTW the answer for part a is supposed to be 63 s and the answer for part b is supposed to be +32m/s... could you show me the appraoch to getting to those answers! thanks

Answer by htmentor(1343)

The height h, is given by:
h = v0*t + 1/2a*t^2
Since v0 = 0, we have h = 1/2a*t^2
t = sqrt(2h/a) = sqrt(2000/0.5) = 63.25 ~ 63 s
For the final velocity use v1 = v0 + at -> v1 = 0.5*63.25 = 31.62 ~ 32 m/s